Problem 3

At the end of 2019, ABC Company’s manufacturing inventory and expense accounts held the following costs.

Work in Process Inventory

Finished Goods Inventory

Cost of Goods Sold

Direct Materials

P120,000

P190,000

P350,000

Direct Labor

P100,000

P250,000

P580,000

Factory Overhead

P60,000

P300,000

P480,000

Alonzo’s accountant applied overhead during the year using a budgeted rate of P8.40 per hour. At year end, they computed the actual rate of P10.00 per machine hour. The beginning balances of both work in process inventory and finished goods inventory were zero.

1. How much is the prorated share of Cost of Goods Sold on the allocated amount of over/under applied overhead?

2. How much is the prorated share of Work in process on the allocated amount of over/under applied overhead?

3. How much is the prorated share of Finished Goods on the allocated amount of over/under applied overhead?
Answer #1
WORKINGS:
1] Factory overhead applied = 60000+300000+480000 = $       840,000
2] Actual hours worked = 840000/8.4 = 100000
3] Actual factory overhead = 100000*10 = $ 1,000,000
4] Overhead underapplied = 1000000-840000 = $       160,000
The above underapplied is to be prorated to WIP,
FG and COGS in the ratio of overhead applied; that is
in the ratio of 60000:300000:480000 or 1:5:8.
ANSWERS:
1] Under applied overhead prorated to COGS = 160000*8/14 = $         91,429
2] Under applied overhead prorated to WIP = 160000*1/14 = $         11,429
3] Under applied overhead prorated to FG = 160000*5/14 = $         57,143
Total under applied overhead prorated $       160,000